OpenStudy (anonymous):

I WILL MEDAL ** Which ordered pair in the form (a, b) is a solution of this equation? 4a – 5b = 24 A. (–2, –5) B. (–5, –2) C. (6, 0) D. (0, 6)

OpenStudy (anonymous):

3x−y4=11 3(4)−44=11 12−44=11 12−1=11 11=11 This means this equation is also correct! Which means "Both are Solutions" There is your Answer!

OpenStudy (anonymous):

So what will be the answer?

OpenStudy (anonymous):

when you simplify & both sides of the equation are equal, u have the answer

OpenStudy (anonymous):

Im in a rush i just need the answer?

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem, just had to show work for you sorry for the wait hahha.

OpenStudy (anonymous):

its fine!